# Robustness of Dark Energy Phenomenology Across Different Parameterizations

### v1

@@ 5.2 Reproducing dark energy phenomenology
Figure 3: Validation of the parametrizations to reproduce the phenomenology of the scalar field. From top to bottom: the hilltop branch of thawing quintessence, the positive mass branch of thawing quintessence, and non-minimal thawing quintessence. Left column: $\chi^{2}$ distribution from the parameterizations fit to the observables computed with the scalar field from Eq. (3.11). The black dashed line corresponds to the degrees of freedom ($\simeq 20$). Right column: 68% and 95% C.L. distributions of the fitted observables at different redshifts (dashed color lines), compared with the distribution obtained with the scalar field (dashed black line). All parametrizations accurately reproduce the theory predictions, but the JBP model does not perform as well as the others.
@@ 5.3 Data constraints and viability of the dark energy models
This is confirmed when we compare the various parametrizations and thawing quintessence in terms of their ability to fit the complete (uncompressed) data. To do so, we find the best fitting model of thawing quintessence (i.e. that maximizes the likelihood ${\cal L}$ or, alternatively, minimizes $\chi^{2}=-2\ln{\hat{\cal L}}$ for the given datasets  [^5] Here, we use the lite version of Planck likelihood to do the minimization as it has substantially fewer parameters than the full likelihood and is thus more efficient.). We do the same with $\Lambda$CDM and with the CPL, JBP, EXP, and BA parameterizations. The best-fit $\chi^{2}$ can can be seen in Table (2) and show that both thawing quintessence and all parametrization can fit the data better than $\Lambda$CDM, in agreement with those of [28]. In particular, while the scalar field improves the $\chi^{2}$ of $\Lambda$CDM only by $\Delta\chi^{2}=-2.9$, the parametrizations obtain $\Delta\chi^{2}=-6.7,\,-6.9,\,-7.1$, for CPL, EXP and BA. The worst performing parametrization is JBP, with $\Delta\chi^{2}=-4.9$, and this worse fit is the reason why in Fig. (4), for JBP, the wedge of the theory projection has a bigger overlap with the data constraints: both JBP and thawing quintessence are relatively close in terms of their ability to fit the expansion history data, meaning that when thawing quintessence is projected into JBP parameter space there will be more overlap than with the other parameterizations which are further away from thawing quintessence in terms of their ability to fit the data.
@@
Table 2: Best fit $\chi^{2}$ values for $\Lambda$CDM and the various dark energy parameterizations and models considered in this paper. In parenthesis, the difference respect to $\Lambda$CDM best fit $\chi^{2}$. These values were obtained by minimizing $\chi^{2}$.
Moving onto non-minimal thawing quintessence, we again find that all parameterizations produce very similar results. However, this time the dark energy models are not confined to a narrow wedge, but rather sweep out across the $(w_{0},w_{a})$ plane for all parametrizations, including covering much of the favored regions picked out by the data as can be seen in Fig. (4). This suggests that this model provides at the very least a comparable fit to the data as the various parameterizations. As discussed in [13], the ability for these models to occupy this region of the $(w_{0},w_{a})$ parameter space is due to their phantom crossing behavior. They have the dynamical freedom to be phantom throughout much of their history, before sharply thawing to cross the phantom divide. However, this thawing can occur in during the redshifts where cosmological probes have robust coverage (see the example models depicted in Fig. (1) here or Fig. (1) in [13]). Consequently, the fitted $(w_{0},w_{a})$ parameters for this kind of dark energy evolution will map these kinds of dark energy models into the lower left parts of the $(w_{0},w_{a})$ plane as this region implies that dark energy is thawing and has undergone significant temporal evolution.
@@ 6 Discussion
The recent DESI results, in conjunction with other cosmological data, paints an intriguing picture regarding the fundamental nature of dark energy. However, as we have seen, the process of interpretting the data involves making what are, to a large extent, arbitrary choices for the parameterization of the dark energy equation of state. In this paper, we explored the robustness of conclusions about dynamical dark energy models—both minimally and non-minimally coupled scalar field models—across four widely used two-parameter families: the Chevallier-Polarski-Linder (CPL), Jassal-Bagla-Padmanabha (JBP), Barboza-Alcaniz (BA), and exponential (EXP) parameterizations [71, 72, 76, 77, 78].
Our analysis reveals that assessing the viability of these dynamical dark energy models is remarkably robust to the choice of parameterization. By mapping the dynamical dark energy models into the $(w_{0},w_{a})$ parameter space for each parameterization, we demonstrated both that our conclusions regarding their viability (or how well they describe the data) are not overly subject to particular choices of parameterization and also that all of these various parameterizations effectively capture the phenomenology of the full dark energy models. This demonstrates that, at least for these very broad and simple kinds of microphysical dark energy models considered here, any of these choices of parameterization are sufficient to interpret the data and make assessments regarding the viability of such dark energy proposals.
@@ Appendix A Compressed data for theory priors
Adopting a particular parametrization allows us, in a simplified way which can be easily visualized, to compare what the data is preferring with where the microphysical model would like to lie. Put another way (and ignoring other cosmological parameters for now), we can compare the likelihood for the parameters $(w_{0},w_{a})$ dictated by the data, $D$, i.e. ${\cal L}(w_{0},w_{a}|D)$, with the priors dictated by a microphysical model with parameters ${\vec{\alpha}}$, i.e. ${\cal P}_{r}(w_{0},w_{a}|{\vec{\alpha}})$. While, ultimately, for any microphysical model, we are interested in the posterior distribution on ${\vec{\alpha}}$, given by
@@
This is exactly analogous to what is done in the study of inflation models. We make observations that constrain the tensor-to-scalar ratio $r$ and the scalar spectral index $n_{s}$, and then use these constraints to determine the observationaly favored region of the $(r,n_{s})$ plane. To evaluate whether or not specific inflationary models are viable, we compute their predictions for $(r,n_{s})$ and compare these to those constraints (e.g. [136, 137, 138, 139, 140]). However, unlike with inflation, as we have seen there are a number of distinct parameterizations for dark energy and these parameters are a bit further from the actual observables. So, in order to actually determine how well a dark energy model maps onto these constraints, we must determine what a dark energy model would predict for $(w_{0},w_{a})$ by using a procedure that mimics what was done for determining where the data itself lies in terms of the parameters. In the dark energy literature, it is common to generate a dark energy model’s equation of state and “fit” this equation of state over some recent range of redshifts using one of the parameterizations directly as a proxy for determining where a model might lie on the $(w_{0},w_{a})$ plane. However, as discussed in [73, 13, 12], because we do not measure the equation of state directly doing this can often lead to highly misleading results. In other words, we must generate the actual observables for a given dark energy model, and then determine what $(w_{0},w_{a})$ parameters best reproduce those observables given the the known errors and uncertainties of the data. This will allow us to make an “apples to apples” comparison.

### v2

@@ 5.2 Reproducing dark energy phenomenology
Figure 3: Validation of the parametrizations to reproduce the phenomenology of the scalar field. From top to bottom: the hilltop branch of thawing quintessence, the positive mass branch of thawing quintessence, and non-minimal thawing quintessence. Left column: $\chi^{2}$ distribution from the parameterizations fit to the observables computed with the scalar field from Eq. (3.12). The black dashed line corresponds to the degrees of freedom ($\simeq 20$). Right column: 68% and 95% C.L. distributions of the fitted observables at different redshifts (dashed color lines), compared with the distribution obtained with the scalar field (dashed black line). All parametrizations accurately reproduce the theory predictions, but the JBP model does not perform quite as well as the others.
@@ 5.3 Data constraints and viability of the dark energy models
This is confirmed when we compare the various parametrizations and thawing quintessence in terms of their ability to fit the complete (uncompressed) data. To do so, we find the best fitting model of thawing quintessence (i.e. that maximizes the likelihood ${\cal L}$ or, alternatively, minimizes $\chi^{2}=-2\ln{\hat{\cal L}}$ for the given datasets [^5] Here, we use the lite version of Planck likelihood to do the minimization as it has substantially fewer parameters than the full likelihood and is thus more efficient.). We do the same with $\Lambda$CDM and with the CPL, JBP, EXP, and BA parameterizations. The best-fit $\chi^{2}$ can can be seen in Table (2) and show that both thawing quintessence and all parametrization can fit the data better than $\Lambda$CDM, in agreement with [29]. In particular, while the scalar field improves the $\chi^{2}$ of $\Lambda$CDM only by $\Delta\chi^{2}=-2.9$, the parametrizations obtain $\Delta\chi^{2}=-6.7,\,-6.9,\,-7.1$, for CPL, EXP and BA. The worst performing parametrization is JBP, with $\Delta\chi^{2}=-4.9$, and this worse fit is the reason why in Fig. (4), for JBP, the wedge of the theory projection has a bigger overlap with the data constraints: both JBP and thawing quintessence are relatively close in terms of their ability to fit the expansion history data, meaning that when thawing quintessence is projected into JBP parameter space there will be more overlap than with the other parameterizations which are further away from thawing quintessence in terms of their ability to fit the data.
@@
Table 2: Best fit $\chi^{2}$ values for $\Lambda$CDM and the various dark energy parameterizations and models considered in this paper. In parenthesis, the difference with respect to $\Lambda$CDM best fit $\chi^{2}$. These values were obtained by minimizing $\chi^{2}$.
Moving onto non-minimal thawing quintessence, we again find that all parameterizations produce very similar results. However, this time the dark energy models are not confined to a narrow wedge, but rather sweep out across the $(w_{0},w_{a})$ plane for all parametrizations, including covering much of the favored regions picked out by the data as can be seen in Fig. (4). This suggests that this model provides at the very least a comparable fit to the data as the various parameterizations. As discussed in [13], the ability for these models to occupy this region of the $(w_{0},w_{a})$ parameter space is due to their phantom crossing behavior. They have the dynamical freedom to be phantom throughout much of their history, before sharply thawing to cross the phantom divide. However, this thawing can occur within the redshifts where cosmological probes have robust coverage (see the example models depicted in Fig. (1) here or Fig. (1) in [13]). Consequently, the fitted $(w_{0},w_{a})$ parameters for this kind of dark energy evolution will map into the lower left parts of the $(w_{0},w_{a})$ plane as this region implies that dark energy is thawing and has undergone significant temporal evolution.
@@ 6 Discussion
The recent DESI results, in conjunction with other cosmological data, paints an intriguing picture regarding the fundamental nature of dark energy. However, as we have seen, the process of interpretting the data involves making what are, to a large extent, arbitrary choices for the parameterization of the dark energy equation of state. In this paper, we explored the robustness of conclusions about dynamical dark energy models—both minimally and non-minimally coupled scalar field models—across four widely used two-parameter models: the Chevallier-Polarski-Linder (CPL), Jassal-Bagla-Padmanabha (JBP), Barboza-Alcaniz (BA), and exponential (EXP) parameterizations [80, 81, 85, 86, 87].
Our analysis reveals that assessing the viability of these dynamical dark energy models is very robust to the choice of parameterization. By mapping the dynamical dark energy models into the $(w_{0},w_{a})$ parameter space for each parameterization, we demonstrated both that our conclusions regarding their viability (or how well they describe the data) are not overly subject to particular choices of parameterization and also that all of these various parameterizations effectively capture the phenomenology of the full dark energy models. This demonstrates that, at least for these very broad and simple kinds of microphysical dark energy models considered here, any of these choices of parameterization are sufficient to interpret the data and make assessments regarding the viability of such dark energy proposals.
@@ Appendix A Compressed data for theory priors
Adopting a particular parametrization allows us, in a simplified way which can be easily visualized, to compare what the data is preferring with where the microphysical model would naturally lie. Put another way (and ignoring other cosmological parameters for now), we can compare the likelihood for the parameters $(w_{0},w_{a})$ dictated by the data, $D$, i.e. ${\cal L}(w_{0},w_{a}|D)$, with the priors dictated by a microphysical model with parameters ${\vec{\alpha}}$, i.e. ${\cal P}_{r}(w_{0},w_{a}|{\vec{\alpha}})$. While, ultimately, for any microphysical model, we are interested in the posterior distribution on ${\vec{\alpha}}$, given by
@@
This is exactly analogous to what is done in the study of inflation models. We make observations that constrain the tensor-to-scalar ratio $r$ and the scalar spectral index $n_{s}$, and then use these constraints to determine the observationaly favored region of the $(r,n_{s})$ plane. To evaluate whether or not specific inflationary models are viable, we compute their predictions for $(r,n_{s})$ and compare these to those constraints (e.g. [147, 148, 149, 150, 151]). However, unlike with inflation, as we have seen there are a number of distinct parameterizations for dark energy and these parameters are a bit further from the actual observables. So, in order to actually determine how well a dark energy model maps onto these constraints, we must determine what a dark energy model would predict for $(w_{0},w_{a})$ by using a procedure that mimics what was done for determining where the data itself lies in terms of the parameters. In the dark energy literature, it is common to generate a dark energy model’s equation of state and “fit” this equation of state over some recent range of redshifts using one of the parameterizations directly as a proxy for determining where a model might lie on the $(w_{0},w_{a})$ plane. However, as discussed in [82, 13, 12], because we do not measure the equation of state directly, doing this can often lead to highly misleading results. In other words, we must generate the actual observables for a given dark energy model, and then determine what $(w_{0},w_{a})$ parameters best reproduce those observables given the the known errors and uncertainties of the data. This will allow us to make an “apples to apples” comparison.